A soccer player, centered with the facing the goal net on a field head-on, needs to kick a soccer ball towards the net from 30 meters away. If the ball is kicked at an incline of 40 degrees from the horizontal (ground), what initial speed in m/s must the soccer player give the ball to score without bouncing before the post? Assume no air resistance.

Report the value as an integer.
Answer #1

x = 30 m

\theta = 40 degree

vo = ?

------------------------------------

we know

x = vo cos \theta * t -------------- (1)

and

yf = yi + vyt + 1/2ayt2

assuming that launch height and landing height of ball is same, we have

0 = vyt + 1/2ay2 -------- (2)

from (1)

t = x / vo cos\theta

put this in (2)

0 = vo sin \theta * ( x / vo cos\theta) - 1/2 * g * (x / vo cos\theta)2

0 = xtan\theta - gx2 / 2vo2 cos2\theta

put the values,

0 = 30 * tan 40 - 9.8 * 302 / 2 * vo2 * cos 2 40

30 * tan 40 = 9.8 * 302 / 2 * vo2 * cos 2 40

vo = 17.3 m/s